40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1449     Entry time: Wed Apr 1 15:47:48 2009
Author: Yoichi 
Type: Update 
Category: Locking 
Subject: 3.8kHz peak looks like a real optical response of the interferometer 
Yoichi, Peter

To see where the 3.8kHz peak comes from, we locked the interferometer with the CARM fed back only to ETM and increased the arm power to 4.
The CARM error signal was taken from the transmission DC (not PO_DC).
The attached plots show the CARM transfer functions taken in this state (called ETM lock in the legends) compared with the ones taken when the CARM is locked by the feedback to the laser frequency (called "Frequency lock").
The first attachment is the TFs from the CARM excitation (i.e. the ETMs were actuated) to the TR_DC and PO_DC signals.

The second attachment is the AO path loop TFs. This is basically the TF from the frequency actuator to the PO_DC error signal.
I injected a signal into the B-excitation channel of the common mode board (with SR785) and measured the TF from TP2B to TP2A of the board.
For the ETM lock case, the AO loop was not closed because I disabled the switch between TP2A and TP1B.

The observation here is that even with no feedback to the laser frequency, the 3.8kHz peak is still present.
This strongly suggests that the peak is a real optical response of the interferometer.

To realize the ETM lock with arm_power=4, I had to tweak the CM loop shape.
I wrote a script to do this (/cvs/cds/caltech/scripts/CM/ETM_CARM_PowerUp).
You can run this script after drstep_bang has finished.
Attachment 1: CARM-ETM-EXC.png  128 kB  | Hide | Hide all
CARM-ETM-EXC.png
Attachment 2: AOpath-TFs.png  93 kB  | Hide | Hide all
AOpath-TFs.png
ELOG V3.1.3-